OpenStudy (anonymous):

f(x) ; g(x)=x+1 ; h(x)=2-x. What is (fogoh)(x) =?! @rizwan_uet can u help me??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and explain and if there's smth i cant understand i'll ask you :)

OpenStudy (anonymous):

wait! where is f(x)??

OpenStudy (anonymous):

3x (i forgot to write)

OpenStudy (anonymous):

f(x)=3x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x)=3x ; g(x)=x+1 ; h(x)=2-x fogoh(x) = f(g(h(x))) = 3((2-x)+1)= 3(2-x+1) = 3(3-x) = 9-3x fogoh(x) = 9-3x ----------> required

OpenStudy (anonymous):

theeeeeeeenks a lot!!!! :)))

OpenStudy (anonymous):

you are most welcomeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

hmm @rizwan_uet also in another question, it's asked fogoh(9) so instead of X i have to use 9,right?

OpenStudy (anonymous):

yes exactly substitute x with 9.

OpenStudy (anonymous):

thanks thanks! and if there's something i can help you.. you're welcome :3